# On the Convergence Analysis of Muon

### v1

@@ Appendix C Star-Convex
C.3 Convergence analysis with 3.1
Theorem C.2.
Under 4.9, 3.1, and 4.10, if we apply Algorithm 2 with $\eta_{t}=\eta$, then
Thus, to reach the precision $f(W_{T})-f^{*}\leq\epsilon$, we can take $\eta=O(\frac{\epsilon}{rLD_{{\rm op}}})$, and the complexity is $O(rLD_{{\rm op}}^{2}\epsilon^{-1}\text{log}\frac{\Delta}{\epsilon})$, where $\Delta=f(W_{0})-f^{*}$.
Moreover, if we apply Algorithm 2 with adaptive learning rates $\eta_{t}=\frac{\|\nabla f(W_{t})\|_{*}}{rL}$, we have
Proof.
Constant stepsize
First, by the $L$ Frobenius norm Lipschitz smoothness of $f$, we have
$\leq f(W_{t})-\eta_{t}\|\nabla f(W_{t})\|_{*}+\frac{rL\eta_{t}^{2}}{2}$ (18)
Then, according to (14), we have
$-\|\nabla f(W_{t})\|_{*}$ $\leq-\frac{f(W_{t})-f^{*}}{D_{{\rm op}}}$
Set $\eta=\eta$ and combine (18) and (14). We have
and
Thus, to reach the precision $f(W_{T})-f^{*}\leq\epsilon$, we can take $\eta=O(\frac{\epsilon}{rLD_{{\rm op}}})$, and the complexity is $O(rLD_{{\rm op}}^{2}\epsilon^{-1}\text{log}\frac{\Delta}{\epsilon})$, where $\Delta=f(W_{0})-f^{*}$.
Adaptive stepsize
$\leq f(W_{t})-\eta_{t}\|\nabla f(W_{t})\|_{*}+\frac{rL\eta_{t}^{2}}{2}$
$=f(W_{t})-\frac{\|\nabla f(W_{t})\|_{*}^{2}}{2rL}.$ (19)
From the star-convex property, we have
Combining Equation 19 and Equation 20, we obtain
Then we have
where we use $\frac{\Delta_{t}}{\Delta_{t-1}}\geq 1$.
After rearranging terms, we obtain
Appendix D Experimental settings
Experiments of Figure 1 and Figure 2 are conducted on NVIDIA H100 GPUs. Experiments of Figure 3 are conducted on Intel(R) Core(TM) i9-14900HX.
In Figure 1 (a), we randomly select a fixed subset of 100 samples from CIFAR-10 and train a fully connected neural network (Table 1) on this subset, serving as the deterministic setting. The learning rates for GD, Adam, Muon (using Newton–Schulz iterations with momentum), Muon_withoutM (without momentum), and Muon_SVD (Algorithm 1) are tuned from the set $\{1e^{-4},5e^{-4},1e^{-3},5e^{-3},1e^{-2},\allowbreak 5e^{-2},1e^{-1}\}$. We conduct experiments with five different random seeds, and report the average results along with one standard deviation in Figure 1 (a). In Figure 1 (b), we train the same neural network using stochastic batches sampled from the entire CIFAR-10 training set, with a batch size of 120. The learning rates for GD, Muon, Muon_withoutM, and Muon_SVD are tuned from $\{1e^{-3},5e^{-3},1e^{-2},5e^{-2},1e^{-1}\}$, while the learning rate for Adam is tuned from $\{1e^{-4},5e^{-4},1e^{-3},5e^{-3}\}$. The model is trained for 5 epochs, and after each epoch, we evaluate the loss on the entire CIFAR-10 training set. The results are presented in Figure 1 (b).
In Figure 2, we randomly select a fixed subset of 60 samples from MNIST. We then train a fully connected neural network (Table 2) with three matrix parameters ($W^{1}\in\mathbb{R}^{128\times 784}$, $W^{2}\in\mathbb{R}^{64\times 128}$, $W^{3}\in\mathbb{R}^{10\times 64}$) on this subset using GD and Muon (Algorithm 2). We first tune the learning rates for GD and Muon from the set $\{1e^{-3},5e^{-3},1e^{-2},5e^{-2},1e^{-1},5e^{-1},1\}$, and find that the optimal learning rate for both is 0.1. Using the selected learning rate, we optimize the neural network with GD and Muon (Algorithm 2), respectively, and record the loss, $J_{t}$, $L_{t}$, $\|\nabla f(W_{t})\|_{*}^{2}$, and $\|\nabla f(W_{t})\|_{\rm F}^{2}$ throughout the optimization process. These quantities are computed with respect to $W^{2}$, i.e., $\nabla f=\nabla_{W^{2}}f(W^{1}_{t},W^{2}_{t},W^{3}_{t})$ and $\nabla^{2}f=\nabla^{2}_{\mathsf{vec}(W^{2})\mathsf{vec}(W^{2})}f_{v}(\mathsf{vec}(W^{1}_{t}),\mathsf{vec}(W^{2}_{t}),\mathsf{vec}(W^{3}_{t}))$, where $f_{v}(\mathsf{vec}(W^{1}_{t}),\mathsf{vec}(W^{2}_{t}),\mathsf{vec}(W^{3}_{t}))=f(W^{1}_{t},W^{2}_{t},W^{3}_{t})$. The results are shown in Figure 2.
@@ Appendix D Experimental settings
In Figure 3(a)(b), we consider the quadratic function $f(W)=\frac{1}{2}Tr((W-W_{\text{opt}})Q(W-W_{\text{opt}})^{\prime})$, $W\in\mathbb{R}^{15\times 20},Q\in\mathbb{R}^{15\times 15}$, where $W_{\text{opt}}$ is randomly chosen and $Q$ is chosen with an ill condition number. Then, we apply GD and Muon to optimize this function, both start from the $W_{0}=0$ with 4000 iterations. We choose the optimal constant stepsize $\frac{1}{L}$ for GD and choose the stepsize for Muon such that Muon can converge in 4000 iterations with the best function value. For each iteration of both algorithms, we record the difference in function value from the optimum and the spectral norm error to the optimal point. The results are shown in Figure 3(a) and (b). In Figure 3(c), we compare the key constant in convergence analysis under quadratic function with ill-conditioned $Q$. We choose $W_{\text{opt}}\sim U(-50,50)$, which means each element of $W_{\text{opt}}$ is i.i.d. random variables drawn from a continuous uniform distribution. Then, we estimate the key constants $D_{{\rm op}}^{2}L_{*}$ and $D_{\rm F}^{2}L$, where we choose the initial point as $W_{0}=0$. We calculate the ratios $\frac{D_{\rm F}^{2}L}{D_{{\rm op}}^{2}L_{*}}$ over $10^{4}$ random samples. The results are shown in Figure 3(c).
D.1 Model Architectures
@@
We present the architectures of the models used in our experiments in the following tables (Table 1, 2). For every Linear module, the bias term is set as false – so the models contain only matrix parameters.
Table 1: Model Architecture for Figure 1
@@
Fully Connected $+$ ReLU $512\times 3072$
Fully Connected $+$ ReLU $256\times 512$
Fully Connected $+$ ReLU $64\times 256$
@@
Table 2: Model Architecture for Figure 2
Layer Type Matrix Shape $(m\times n)$
Fully Connected $+$ ReLU $128\times 784$
Fully Connected $+$ ReLU $64\times 128$
Fully Connected $10\times 64$
Appendix E Muon with Newton-Schulz iterations
As discussed in Section 3.1, computing SVD is usually expensive in practice, especially when the dimensions are large. Therefore, a common implementation of Muon employs the Newton-Schulz iterations Bernstein and Newhouse [2024], Jordan et al. [2024] to approximate the orthogonalization process (Algorithm 3). In the original paper of Muon [Jordan et al., 2024], they define their Newton-Schulz iterations (line 6 in Algorithm 3) in the following way. First, the input $M_{t}$ will be normalized as $X_{0}=M_{t}/\|M_{t}\|_{\rm F}$. Then, $X_{k}$ is computed from $X_{k-1}$ according to the following update rule:
$X_{k}=aX_{k-1}+b(X_{k-1}X_{k-1}^{\top})X_{k-1}+c(X_{k-1}X_{k-1}^{\top})^{2}X_{k-1}.$
Following the analysis in Jordan et al. [2024] and supposing the SVD of $M_{t}$ is $U_{t}S_{t}V_{t}^{\top}$, we have $X_{0}=U_{t}\bar{S}_{t}V_{t}^{\top}$, where $\bar{S}_{t}=S_{t}/\|S_{t}\|_{\rm F}$. We can note that
$X_{1}=$ $aX_{0}+b(X_{0}X_{0}^{\top})X_{0}+c(X_{0}X_{0}^{\top})^{2}X_{0}$
$=$ $(aI+bU\bar{S}^{2}U^{\top}+cU\bar{S}^{4}U^{\top})U\bar{S}V^{\top}$
$=$ $U(a\bar{S}+b\bar{S}^{3}+c\bar{S}^{5})V^{\top}.$
Define $\varphi(x)=ax+bx^{3}+cx^{5}$. Then after $k$ iterations, the output is $O_{t}=X_{k}=U_{t}\varphi^{k}(\bar{S}_{t})V_{t}^{\top}$. In Jordan et al. [2024], they set $k=5$, $a=3.4445$, $b=-4.7750$, and $c=2.0315$. Note that $\bar{S}_{t}=S_{t}/\|S_{t}\|_{\rm F}$ is normalized; thus, its singular values are in $[0,1]$. We can plot $\varphi^{5}(x)$ in $[0,1]$ and get Figure 4. Note that there is a steep jump around $x=0$ in Figure 4 and if we define a threshold $\varepsilon$ such that for $x\in[\varepsilon,1]$, $\varphi^{5}(x)\in[0.65,1.25]$, we can have that this threshold $\varepsilon$ is very small and most function values of $\varphi^{5}(x)$ lie in $[0.65,1.25]$. Thus, $O_{t}=U_{t}\varphi^{k}(\bar{S}_{t})V_{t}^{\top}$ serves as a low-cost yet effective approximation to $U_{t}V_{t}^{\top}$.
Algorithm 3 Muon with Newton–Schulz iterations
2:for $t=0,1,\ldots,T-1$ do
3:     Sample batch $\{\xi_{t,i}\}_{i=1}^{B}$ uniformly
5:     If $t>0$, $M_{t}=\beta M_{t-1}+(1-\beta)G_{t}$. If $t=0$, $M_{0}=G_{0}$.
6:     $O_{t}=\text{NewtonSchulz}(M_{t})$
7:     $W_{t+1}=W_{t}-\eta_{t}O_{t}$
8:end for
Figure 4: $\varphi^{5}(x)$ with $\varphi(x)=ax+bx^{3}+cx^{5}$ and $a=3.4445$, $b=-4.7750$, $c=2.0315$. Similar to the Figure 4 in Jordan et al. [2024]. Line 0.65 and 1.25 are just for illustrative purposes; one can actually choose tighter bounds.

### v2

@@ Appendix C Star-Convex
Appendix D Low-rank Structure of Real-World Data Matrices
(a) Spectra of $X_{\text{CIFAR10}}\in\mathbb{R}^{3072\times 1000}$
(b) Spectra of $X_{\text{Gaussian, 1}}\in\mathbb{R}^{3072\times 1000}$
(c) Spectra of $X_{\text{Text}}\in\mathbb{R}^{768\times 836}$
(d) Spectra of $X_{\text{Gaussian, 2}}\in\mathbb{R}^{768\times 836}$
Figure 5: Spectra of $X_{\text{CIFAR10}}$, $X_{\text{Text}}$, $X_{\text{Gaussian, 1}}$ and $X_{\text{Gaussian, 2}}$.
In addition to demonstrating the low-rank property of MNIST in Figure 2, we also test and verify similar low-rank structures on the CIFAR-10 [Krizhevsky et al., 2009] and Shakespeare datasets [^2] https://raw.githubusercontent.com/karpathy/char-rnn/master/data/tinyshakespeare/input.txt.
Following a procedure similar to that in Section 4.2.1, for CIFAR-10 we randomly sample 1000 samples from training dataset and construct the matrix $X_{\text{CIFAR10}}\in\mathbb{R}^{3072\times 1000}$. We then compute its singular values and compare them with those of a Gaussian random matrix of the same size, $X_{\text{Gaussian, 1}}\in\mathbb{R}^{3072\times 1000}$. The results are shown in Figure 5(a)(b), where we observe that the singular values of $X_{\text{CIFAR10}}$ are significantly more concentrated than $X_{\text{Gaussian, 1}}$.
For the Shakespeare dataset, we take the first 3000 characters and use the RoBERTa [Liu et al., 2019] tokenizer and embedding model to convert the text into an embedding matrix $X_{\text{Text}}\in\mathbb{R}^{768\times 836}$, where 768 is the embedding dimension and 836 is the token length. We compute its singular values and compare them with those of a Gaussian random matrix of the same size, $X_{\text{Gaussian, 2}}\in\mathbb{R}^{768\times 836}$. As shown in Figure 5(c)(d), similar to the image datasets (MNIST and CIFAR-10), the text embedding matrix also exhibits a pronounced low-rank structure.
Appendix E Experimental Settings
Experiments of Figure 2, Figure 3, and Figure 4 are conducted on NVIDIA RTX 6000 Ada GPUs. Experiments of Figure 1 are conducted on Intel(R) Core(TM) i9-14900HX.
@@ Appendix E Experimental Settings
In Figure 1, we consider the quadratic function $f(W)=\frac{1}{2}\text{tr}\left((W-W_{\text{opt}})^{\top}Q(W-W_{\text{opt}})\right)$, $W\in\mathbb{R}^{15\times 20},Q\in\mathbb{R}^{15\times 15}$, where $W_{\text{opt}}$ is randomly chosen and $Q$ is chosen with an ill condition number. Then, we apply GD and Muon to optimize this function, both start from the $W_{0}=0$ with 4000 iterations. We choose the optimal constant stepsize $\frac{1}{L}$ for GD and choose the stepsize for Muon such that Muon can converge in 4000 iterations with the best function value. For each iteration of both algorithms, we record the difference in function value from the optimum and the spectral norm error to the optimal point. The results are shown in Figure 1(a) and (b). In Figure 1(c), we compare the key constant in convergence analysis under quadratic function with ill-conditioned $Q$. We choose $W_{\text{opt}}\sim U(-50,50)$, which means each element of $W_{\text{opt}}$ is i.i.d. random variables drawn from a continuous uniform distribution. Then, we estimate the key constants $D_{{\rm op}}^{2}L_{*}$ and $D_{\rm F}^{2}L$, where we choose the initial point as $W_{0}=0$. We calculate the ratios $\frac{D_{\rm F}^{2}L}{D_{{\rm op}}^{2}L_{*}}$ over $10^{4}$ random samples. The results are shown in Figure 1(c).
In Figure 3, we randomly select a fixed subset of 120 samples from MNIST. We then train a fully connected neural network (Table 1) with three matrix parameters ($W^{1}\in\mathbb{R}^{128\times 784}$, $W^{2}\in\mathbb{R}^{64\times 128}$, $W^{3}\in\mathbb{R}^{10\times 64}$) on this subset using GD and Muon (Algorithm 2). We tuned the learning rates for GD and Muon to their optimal values. During the optimization procedure, we record the loss, $J_{t}$, $L_{t}$, $\|\nabla f(W_{t})\|_{*}^{2}$, and $\|\nabla f(W_{t})\|_{\rm F}^{2}$. These quantities are computed with respect to $W^{2}$, i.e., $\nabla f=\nabla_{W^{2}}f(W^{1}_{t},W^{2}_{t},W^{3}_{t})$ and $\nabla^{2}f=\nabla^{2}_{\overline{\mathrm{vec}}(W^{2})\overline{\mathrm{vec}}(W^{2})}f_{v}(\overline{\mathrm{vec}}(W^{1}_{t}),\overline{\mathrm{vec}}(W^{2}_{t}),\overline{\mathrm{vec}}(W^{3}_{t}))$, where $f_{v}(\overline{\mathrm{vec}}(W^{1}_{t}),\overline{\mathrm{vec}}(W^{2}_{t}),\overline{\mathrm{vec}}(W^{3}_{t}))=f(W^{1}_{t},W^{2}_{t},W^{3}_{t})$.
@@
Table 1: MLP model used in Figure 3. For every Linear module, the bias term is set as false, so this model contain only matrix parameters.
@@
Fully Connected ++ ReLU $128\times 784$
Fully Connected ++ ReLU $64\times 128$
@@
In Figure 4, we adopted the code from nanoGPT [^3] https://github.com/karpathy/nanoGPT and trained a 6-layer GPT-2–style model on the Shakespeare character dataset. Model architectural and training hyperparameters can be found in Table 2. We trained the model using AdamW and Muon, respectively. When using Muon, we applied the tricks proposed in Liu et al. [2025a], i.e., using weight decay and adjusting the per-parameter update scale. Muon was used only to optimize the matrix parameters in the non-embedding layers, while the embedding layers and vector parameters were optimized with AdamW.
Table 2: Model architectural and training hyperparameters for the experiments in Figure 4.
Hyperparameter Value
Number of Layers 6
Number of Heads 6
Hidden Dimension 384
Dropout 0
Batch Size 64
Training Steps 1000